OpenStudy (anonymous):

The ratio of the number of video games to the number of movies in the library is 3 : 9. After 16 movies were borrowed from the library, the ratio of games to movies became 3 : 6. How many movies were left after 16 were borrowed? 6 movies 24 movies 32 movies 64 movies

OpenStudy (anonymous):

help

OpenStudy (jack1):

video games to movies= 3/9 = 1/3 - 16 movies new ratio = 3/6 = 2/3 so 16 video games and 48 movies total ratio = 1/3 minus 16 movies = 16 video games and 32 movies ratio = 2/3 fits, answer is C... understand how, yeah?